# A scalable adaptive deep Koopman predictive controller for real-time optimization of mixed traffic flow

### v1

@@ V-B Baseline models and experiment settings
LSTM utilizes LSTM-based encoder-decoder architecture to predict vehicle trajectories [51].
Koopnet is an ablation version of AdapKoopnet where the extraction of driving characteristics semantic information from the trajectories context of vehicle is eliminated.
N-AdapKoopnet is a variant of AdapKoopnet where the decoder is replaced with fully connected layers featuring non-linear activation functions.
N-Koopnet features an encoder composed of fully connected layers with non-linear activation functions.
S-AdapKoopnet employs a modified architecture in which the width of each layer is halved compared to AdapKoopnet while retaining the same structure. This modification aims to reduce the computational load, especially when implementing optimization for large-scale mixed traffic flow.
PyTorch 2.1.0 framework is utilized to construct the predictive model, and end-to-end training was performed under a platform with Intel Core i9-13900K processor and NVIDIA RTX 4090 GPU. The processed car-following trajectories were split into training, validation, and test sets in a ratio of 7:1:2. The model relevant hyperparameters are detailed in Table I.
@@
V-C Trajectory prediction experiment results
V-C1 HDVs trajectory prediction performance comparison
Table II statistics the performance index results of AdapKoopNet and baseline models for trajectory prediction under different prediction horizons. In terms of velocity, when the prediction step is 5 (0.6 seconds), the RMSE of LSTM is the smallest at 0.084m. The reason may be due to the architectural design of the LSTM encoder-decoder and the excellent gate control unit structure, which has advantages in short time domain prediction. When the prediction step increases to 1.2 seconds, AdapKoopNet has a significant advantage, with an RMSE of 0.145, which means it has better long-term prediction ability. AdapKoopnet considers the multi-step evolution loss of linear space during the training process, giving up some short-term prediction performance in exchange for the average prediction performance in the entire prediction time domain. Although Koopman theory provides a global linear expression of a dynamic system in an infinite-dimensional space, AdapKoopnet ultimately implements predictions in a finite-dimensional (128-dimensional) space, which is bound to be affected by linear systems, resulting in some losses. Simply put, with the help of the powerful feature extraction capability of the attention mechanism and the adaptive driving feature extraction architecture based on driving scenarios, the AdapKoopnet series models have achieved performance comparable to LSTM in terms of short-term prediction performance, and better than it in terms of long-term prediction performance.
As the prediction step further increases, N-AdapKoopNet outperforms AdapKoopNet in performance, reflecting the loss of prediction performance caused by the adoption of linear decoders, the average performance at each prediction step size is still the most outstanding for AdapKoopNet. Subsequently, comparing the distance between the front of the vehicle, the most effective models are LSTM and N-AdapKoopNet. However, the decoder of this model is non-linear and can indeed achieve good trajectory prediction results. However, it is also not suitable for subsequent CAVs prediction control. Taking into account the accuracy of trajectory prediction and the feasibility of subsequent control for real-time traffic flow optimization, it is evident that AdapKoopNet has significant advantages. In addition, in order to balance prediction performance and inference cost, the large-scale traffic flow simulation experiment in Section 6.3 is based on S-AdapKoopnet.
@@ V-C1 HDVs trajectory prediction performance comparison
Fig. 6 (a) shows the indicator line graph of AdapKoopNet and the baseline models under complete 15 prediction steps, where the red line represents AdapKoopNet. The comparison results mainly including the following two aspects. Firstly, regardless of the prediction range, the AdapKoopNet model outperforms KoopNet, indicating that it effectively and accurately captures the underlying data features and achieving personalized driving behavior prediction. Secondly, as the prediction range increases, although the prediction performance of AdapKoopNet has declined, it still maintains accuracy and feasibility. The comparative analysis with the baseline models is similar to Table II and will not be further elaborated.
Fig. 6 (b) focuses specifically on the prediction step scene of 1.8s. 300 batches are randomly selected from the test set, and the predicted and true values of headway and velocity are visualized, with the color axis representing the probability density. It can be observed that the velocity of most trajectory appears between 20 and 30 meters, while the headway remains around 30 to 50 meters. In addition, the vast majority of trajectory samples are attached near the diagonal, which means that the predicted values are very close to the true values, reflecting the strong ability of AdapKoopNet to capture trajectory evolution features and make accurate predictions. Of course, there are some samples with significant deviations, but the density of these samples is extremely low and almost does not affect the overall trajectory prediction performance. Besides, benefiting from the rolling optimization method adopted in the AdapKoopPC, the negative impact of these deviation samples can be ignored.
V-C2 Potential scenarios recognition and clustering results
@@
The first thing to explore is what all trajectory samples output after passing through the multi-head driving scenario recognition module. Fig. 7 shows the distribution of trajectory samples for various driving scenarios learned and clustered. Fig. Fig. 7 (a), (b) and (c)show the headway-velocity relationship of trajectory samples under three driving scenarios. The color axis represents the proportion of samples, with the proportion of trajectory samples increasing as the color approaches red. It is obvious that Fig. 7 (a) tends towards medium to high velocity driving scenario, with the highest proportion of sample clusters distributed at velocity of around 25m/s, accompanied by a moderate headway of 30 meters to 60 meters. The distribution of trajectory samples in Fig. 7 (b) is relatively uniform, including scenario with small headway at low velocity and large headway at high velocity. Fig. 7 (c) corresponds to the third driving scenario, where the overall trajectory exhibits low velocity accompanied by small headway. Fig. 7 (d), (e), and (f) depicts the relationship between the average velocity difference, average headway, and average velocity of trajectory samples in three driving scenarios, with the color axis representing velocity. Significantly, all trajectory samples are clearly clustered into three driving scenarios, each with similar repetitive patterns. Of course, the characteristics of all trajectories in each posture scenario are not entirely the same. Overall, AdapKoopNet can adaptively extract features from a large number of trajectory samples and cluster them, ultimately forming these three potential driving scenarios.
V-C3 Scenario inherent temporal correlation and the driving characteristics dynamic temporal correlation
@@
Fig. 8 reveals this specific driving scenario inherent temporal correlation and the driving characteristics dynamic temporal correlation. 8 (a) corresponds to the output of the driving characteristics semantic module shown in 3 (b) in AdapKoopNet. It is worth noting that as mentioned earlier, AdapKoopNet adaptively learns the feature weights of each trajectory sample belonging to three different scenarios, and then clusters them into potential scenarios according to their propensity. Therefore, more precisely, what is extracted here is the inherent temporal correlation corresponding to each driving scenario. Specifically, driving scenario 1 and driving scenario 3 are quite similar, both reflecting an overall trend of lower temporal correlation as the historical time step increases. The difference lies in that the former is not as extreme as the latter, and Scenario 3 strongly relies on the nearest time step and has almost no correlation on distant historical trajectories. There is a significant difference between driving scenario 2 and the above two driving scenarios, characterized by a more stable temporal correlation on the historical trajectory of each time step, rather than being more affected as time approaches.
@@
V-C4 Validation of scenarios recognition effectiveness
@@
VI Experiment Part II: AdapKoopPC for mixed traffic flow
This section aims to comprehensively evaluate the effectiveness and superiority of the proposed AdapKoopPC in mitigating traffic oscillations in mixed traffic flows. Specifically, it covers small-scale experiments and large-scale experiments.
VI-A Simulation settings
The small-scale simulated mixed traffic flow consists of 2 CAVs, 2 HDVs(Truck) and 6 HDVs(Car). The large-scale simulated mixed traffic flow consists of 50 vehicles, of which CAVs and HDVs(Car) account for 80%, and HDVs(Truck) account for 20%. A certain proportion of trucks is set up to build a heterogeneous simulation environment, with the purpose of proving the modeling and prediction capabilities of AdapKoopPC for heterogeneous HDVs. IDM is used as the simulation control model of HDVs, and its parameters are calibrated and obtained from the naturalistic driving dataset using genetic algorithm. The simulation is conducted on an open road, and the head vehicle of the mixed traffic flow evolves according to the following equation to simulate traffic oscillations [52]:
Except for the first vehicle, which must be a CAV, the arrangement of vehicles in the local traffic flow is randomly generated based on the proportion of vehicle types. The control input for CAVs is solved using the SLSQP optimizer from the Scipy library, and the simulation is conducted on a laptop equipped with the Apple M2 chip. The simulation experiment parameters are shown in Table III.
| Parameter | Value | Parameter | Value | |
| 1.13 | 1.5 | |||
| 4 | 4 | |||
| IDM | 8.16 | 9.66 | ||
| 1.13 | 1.72 | |||
| 35.96 | 54.25 | |||
| 4.24 | 11.82 | |||
| 20 | 0 | |||
| 150 | 150 | |||
| AdapKoopPC | -6 | 6 | ||
| 20 | 10 | |||
| 6 | -6 | |||
| 10 | 2 | |||
| Simulation | Duration(s) | 180 | Interval(s) | 0.12 |
VI-B Real-time control optimization experiment for small-scale mixed traffic flow
@@
LTI-MPC: On the premise that the car-following model adopted by HDVs in the mixed traffic flow is known, linear time-invariant MPC is used to optimize the mixed traffic flow based on the linearized car-following model [18].
Deep-LCC: A data-driven non-parametric predictive control framework for mitigating traffic flow velocity oscillations. And the evolution dynamics of mixed traffic flow are learned online using the pre-collected trajectories [18].
@@
VI-B1 Mixed traffic flow evolution results
@@
Fig. 10 and Fig. 11 depict and compares the headway and acceleration evolution of small-scale mixed traffic flow under different control methods. The yellow dashed line represents the head vehicle, while the blue and orange thick solid lines represent the two CAVs under various control methods. The remaining 8 vehicles are all HDVs, with HDV8 and HDV10 being trucks and the rest being cars. Fig. 10 (a) and Fig. 11 (a) depict the evolution of the mixed traffic flow corresponding to no control, where the CAV solely considers optimizing its following trajectory with respect to the head vehicle, without taking into account its own guiding influence on the subsequent HDVs. After the head vehicle experiences disturbances starting at 4.8s and begins to oscillate, and the entire traffic flow exhibits substantial acceleration and deceleration. The evolution of the blue line in the middle subgraph indicates that as the CAV decelerates with the leading vehicle, the velocity of the entire mixed traffic flow also begins to oscillate significantly. The following HDVs exhibit significant fluctuations in their headways. Taking the first HDV following behind as an example, the headway drops to 33 meters and then continues to rise to around 72 meters, indicating poor stability due to the promotion of disturbances through the entire traffic flow.
@@
Fig. 10 (b) and Fig. 11 (b) shows the mixed traffic flow evolution after implementing the LTI-MPC. There is a change that needs to be explained in advance, which is that the first CAV did not maintain a constant velocity with the leading vehicle at the beginning. This is because LTI-MPC imposes a penalty on the error between the actual headway and the expected headway. Once there is a small error between the initial headway set in the experiment and the expected headway after traffic flow operates, it will lead to fluctuations in the initial stage of mixed traffic flow evolution, which is difficult to be avoided in the experimental setup. It is obvious that the amplitude of the velocity oscillation of the first CAV also begins to oscillate after receiving the oscillation propagation from the leading vehicle, and the following HDV also oscillates with it. The overall oscillation amplitude decreases relative to Basic-MPC. In terms of headway, taking the second HDV indicated by the red line as an example, the oscillation amplitude of the headway improved due to the influence of the CAV ahead, reaching a maximum front distance of about 50 meters at 40 seconds. Subsequently, the second CAV also optimized its driving behavior under the action of LTI-MPC, and the trajectories of the following two trucks also fluctuated accordingly, maintaining a distance near the expected headway.
@@
The evolution and improvement of the mixed traffic flow under the Deep-LCC control scenario are shown in Fig. 10 (c) and Fig. 11 (c). Similarly, the initial evolution stage, like the LTI-MPC, requires the first CAV to accelerate in order to achieve equilibrium headway. We can clearly observe that there is little difference between DeeP-LCC and LTI-MPC in overall traffic flow evolution improvement, achieving performance comparable to LTI-MPC. The specific details will not be repeated in the description. Of course, Deep-LCC does not require prior knowledge of flow dynamics, and its advantage lies in its data-driven non parametric strategy.
@@
Overall, these comparative results further demonstrate that AdapKoopPC can effectively and proactively guide the subsequent HDVs and maintains the smooth driving of the mixed traffic flow. More importantly, whether it is LTI-MPC or Deep-LCC, it is necessary to utilize a given model or simulation to generate some trajectories. However, these trajectories are highly dependent on the expected headway, which has limitations. AdapKoopPC adopts a penalty of velocity difference, which matches the expected goal of most drivers of vehicles in a following state. This will not cause discomfort to the driver and is more conducive to optimize the control of the mixed traffic flow.
VI-B2 Computation time comparison
Table IV states the real-time computing time for different control methods at each time step, with a time step of 0.12 seconds. The average computation time for LTI-MPC, Deep LCC, and AdapKoopPC is 0.004s, 0.326s, and 0.009s, respectively. To investigate its reasons, LTI-MPC does not require neural network encoding, and the mixed traffic flow state dimension is low, so the solution is the fastest. As for Deep-LCC, it was shown to be 0.05s in theprevious work. The online computation time of Deep-LCC has shown significant fluctuations in our experiments. When simulating mixed traffic flow scenarios similar to pre-generated trajectory scenarios, the average computation delay is 0.326 seconds. When there is a significant difference between the two scenarios, the model training solution needs to iterate to the maximum number of manually set iterations. Therefore, the delay depends on this value and will not be lower than 0.326 seconds, which may be due to a lack of generalization ability. After the impact of neural network encoding time and significant increase in state dimension, the online computation of AdapKoopPC yields an average time of 0.009 seconds, and the sampling interval is 0.12 seconds, which demonstrates the feasibility of implementing real-time mixed traffic flow control. It is worth mentioning that the simulation experiments of LTI-MPC and Deep-LCC are based on the MATLAB quadprog optimizer, which has better solving performance compared to the SciPy adopted in AdapKoopPC. Therefore, the online computation time of AdapKoopPC can still be further improved.

### v3

@@ 5.2 Baseline models and experiment settings
LSTM utilizes LSTM-based encoder-decoder architecture to predict vehicle trajectories (Park et al., 2018).
Koopnet is an ablation version of AdapKoopnet where the extraction of driving characteristics semantic information from the trajectories context of vehicle is eliminated. This emphasizes the importance of the driving characteristic extraction block.
N-AdapKoopnet is a variant of AdapKoopnet where the decoder is replaced with fully connected layers featuring non-linear activation functions. This demonstrates the impact of linear decoder on the prediction results.
N-Koopnet is similar to N-AdapKoopnet, features an encoder composed of fully connected layers with non-linear activation functions.
S-AdapKoopnet employs a modified architecture in which the width of each layer is halved compared to AdapKoopnet while retaining the same structure. This modification aims to reduce the computational load, especially when implementing optimization for large-scale mixed traffic systems.
PyTorch 2.1.0 framework is utilized to construct the predictive model, and end-to-end training was performed under a platform with Intel Core i9-13900K processor and NVIDIA RTX 4090 GPU. The processed car-following trajectories were split into training, validation, and test sets in a ratio of 7:1:2. The model relevant hyperparameters are detailed in 1.
@@
5.3 Trajectory prediction experiment results
5.3.1 HDVs trajectory prediction performance comparison
Table 2 statistics the performance index results of AdapKoopNet and other baseline models for trajectory prediction under different prediction horizons, including the RMSE of velocity and headway. In terms of velocity, when the prediction step is 5 (0.6 seconds), the RMSE of LSTM is the smallest at 0.084m. The reason may be due to the architectural design of the LSTM encoder-decoder and the excellent gate control unit structure, which has advantages in short time domain prediction. When the prediction step increases to 1.2 seconds, AdapKoopNet has a significant advantage, with an RMSE of 0.145m, which means it has better long-term prediction ability. AdapKoopnet considers the multi-step evolution loss of linear space during the training process, giving up some short-term prediction performance in exchange for the average prediction performance in the entire prediction time domain. It is also worth mentioning that although Koopman theory provides a global linear expression of a dynamic system in an infinite-dimensional space, AdapKoopnet ultimately implements predictions in a finite-dimensional (128-dimensional) space, which is bound to be affected by linear systems, resulting in some losses. Simply put, with the help of the powerful feature extraction capability of the attention mechanism and the adaptive driving feature extraction architecture based on driving scenarios, the AdapKoopnet series models have achieved performance comparable to LSTM in terms of short-term prediction performance, and better than it in terms of long-term prediction performance.
Although as the prediction step further increases, N-AdapKoopNet outperforms AdapKoopNet in performance, reflecting the loss of prediction performance caused by the adoption of linear decoders, the average performance at each prediction step size is still the most outstanding for AdapKoopNet. Subsequently, comparing the distance between the front of the vehicle, the most effective models are LSTM and N-AdapKoopNet. However, the decoder of this model is non-linear and can indeed achieve good trajectory prediction results. However, it is also not suitable for subsequent CAVs prediction control. Taking into account the accuracy of trajectory prediction and the feasibility of subsequent control for real-time traffic flow optimization, it is evident that AdapKoopNet has significant advantages. In addition, in order to balance prediction performance and inference cost, the large-scale hybrid transportation system simulation experiment in Section 6.3 is based on S-AdapKoopnet.
@@ 5.3.1 HDVs trajectory prediction performance comparison
Fig. 6 (a) shows the indicator line graph of AdapKoopNet and the baseline models under complete 15 prediction steps, where the red line represents AdapKoopNet. The comparison results are obvious, mainly including the following two aspects. Firstly, regardless of the prediction range, the AdapKoopNet model outperforms KoopNet, indicating that it effectively and accurately captures the underlying data features and achieving personalized driving behavior prediction. Secondly, as the prediction range increases, although the prediction performance of AdapKoopNet has declined, it still maintains accuracy and feasibility. The comparative analysis with the baseline models is similar to Table 2 and will not be further elaborated.
Fig. 6 (b) focuses specifically on the prediction step scene of 1.8s. 300 batches are randomly selected from the test set, and the predicted and true values of headway and velocity are visualized, with the color axis representing the probability density. It can be observed that the velocity of most trajectory appears between 20 and 30 meters, while the headway remains around 30 to 50 meters. In addition, the vast majority of trajectory samples are attached near the diagonal, which means that the predicted values are very close to the true values, reflecting the strong ability of AdapKoopNet to capture trajectory evolution features and make accurate predictions. Of course, there are some samples with significant deviations, but it can be seen that these scattered points are all dark blue, indicating that the density of these samples is extremely low and almost does not affect the overall trajectory prediction performance. Of course, benefiting from the rolling optimization method adopted in the AdapKoopPC, the negative impact of these deviation samples on the subsequent mixed traffic flow optimization control can be ignored.
5.3.2 Potential scenarios recognition and clustering results
@@
The first thing to explore is what all trajectory samples output after passing through the multi-head driving scenario recognition module. Fig. 7 shows the distribution of trajectory samples for various driving scenarios learned and clustered. Fig. 7 (a), (b) and (c)show the headway-velocity relationship of trajectory samples under three driving scenarios. The color axis represents the proportion of samples, with the proportion of trajectory samples increasing as the color approaches red. It is obvious that Fig. 7 (a) tends towards medium to high velocity driving scenario, with the highest proportion of sample clusters distributed at velocity of around 25m/s, accompanied by a moderate headway of 30 meters to 60 meters. The distribution of trajectory samples in Fig. 7 (b) is relatively uniform, including scenario with small headway at low velocity and large headway at high velocity. Fig. 7 (c) corresponds to the third driving scenario, where the overall trajectory exhibits low velocity accompanied by small headway. Fig. 7 (d), (e), and (f) depicts the relationship between the average velocity difference, average headway, and average velocity of trajectory samples in three driving scenarios, with the color axis representing velocity. Significantly, all trajectory samples are clearly clustered into three driving scenarios, each with similar repetitive patterns. Of course, the characteristics of all trajectories in each posture scenario are not entirely the same. Overall, AdapKoopNet has the ability to adaptively extract features from a large number of trajectory samples and cluster them, ultimately forming these three potential driving scenarios.
5.3.3 Scenario inherent temporal correlation and the driving characteristics dynamic temporal correlation
@@
Fig. 8 reveals this specific driving scenario inherent temporal correlation and the driving characteristics dynamic temporal correlation. Fig. 8 (a) corresponds to the output of the driving characteristics semantic module shown in 3 (b) in AdapKoopNet. It is worth noting that as mentioned earlier, AdapKoopNet adaptively learns the feature weights of each trajectory sample belonging to three different scenarios, and then clusters them into potential scenarios according to their propensity. Therefore, more precisely, what is extracted here is the inherent temporal correlation corresponding to each driving scenario. Specifically, driving scenario 1 and driving scenario 3 are quite similar, both reflecting an overall trend of lower temporal correlation as the historical time step increases. The difference lies in that the former is not as extreme as the latter, and Scenario 3 strongly relies on the nearest time step and has almost no correlation on distant historical trajectories. There is a significant difference between driving scenario 2 and the above two driving scenarios, characterized by a more stable temporal correlation on the historical trajectory of each time step, rather than being more affected as time approaches.
@@
5.3.4 Validation of scenarios recognition effectiveness
@@
6 Experiment Part II: AdapKoopPC for mixed traffic system
This section aims to comprehensively evaluate the effectiveness and superiority of the proposed AdapKoopPC in mitigating traffic oscillations in mixed traffic flows. Specifically, it covers small-scale experiments and large-scale experiments. The former aims to compare the spatio-temporal evolution trajectory of each vehicle in the mixed traffic system based on AdapKoopPC and based on other benchmark control methods, see Section 6.2; the latter is to verify the application of AdapKoopPC for the performance and generalization ability of large-scale mixed transportation systems, see Section 6.3.
6.1 Simulation settings
The small-scale simulated mixed traffic system consists of 2 CAVs, 2 HDVs(Truck) and 6 HDVs(Car). The large-scale simulated mixed traffic system consists of 50 vehicles, of which CAVs and HDVs(Car) account for 80%, and HDVs(Truck) account for 20%. A certain proportion of trucks is set up to build a heterogeneous simulation environment, with the purpose of proving the modeling and prediction capabilities of AdapKoopPC for heterogeneous HDVs. IDM is used as the simulation control model of HDVs, and its parameters are calibrated and obtained from the naturalistic driving dataset using genetic algorithm. The simulation is conducted on an open road, and the head vehicle of the mixed traffic system evolves according to the following equation to simulate traffic oscillations (Ruan et al., 2022):
Except for the first vehicle, which must be a CAV, the arrangement of vehicles in the system is randomly generated based on the proportion of vehicle types. The control input for CAVs is solved using the SLSQP optimizer from the Scipy library, and the simulation is conducted on a laptop equipped with the Apple M2 chip. The simulation experiment parameters are shown in Table 3.
| Parameter | Value | Parameter | Value | Parameter | Value | |
| 1.13 | 1.5 | 35.96 | ||||
| IDM | 4 | 4 | 54.25 | |||
| 8.16 | 9.66 | 4.24 | ||||
| 1.13 | 1.72 | 11.82 | ||||
| 20 | 0 | 6 | ||||
| AdapKoopPC | 150 | 150 | -6 | |||
| -6 | 6 | 10 | ||||
| 20 | 10 | 2 | ||||
| Simulation | Duration(s) | 180 | Interval(s) | 0.12 | Truck proportion | 20% |
6.2 Real-time control optimization experiment for small-scale mixed traffic system
@@
LTI-MPC: On the premise that the car-following model adopted by HDVs in the mixed traffic system is known, linear time-invariant MPC is used to optimize the mixed traffic system based on the linearized car-following model (Wang et al., 2023a).
Deep-LCC: A data-driven non-parametric predictive control framework for mitigating traffic flow velocity oscillations. And the evolution dynamics of mixed traffic systems are learned online using the pre-collected trajectories (Wang et al., 2023a).
@@
6.2.1 Mixed traffic system evolution results
@@
Fig. 10 and Fig. 11 depict and compares the headway and acceleration evolution of small-scale mixed traffic system under different control methods. The yellow dashed line represents the head vehicle, while the blue and orange thick solid lines represent the two CAVs under various control methods. The remaining 8 vehicles are all HDVs, with HDV8 and HDV10 being trucks and the rest being cars. Fig. 10 (a) and Fig. 11 (a) depict the evolution of the mixed traffic system corresponding to no control, where the CAV solely considers optimizing its following trajectory with respect to the head vehicle, without taking into account its own guiding influence on the subsequent HDVs. After the head vehicle experiences disturbances starting at 4.8s and begins to oscillate, and the entire system exhibits substantial acceleration and deceleration. The evolution of the blue line in the middle subgraph indicates that as the CAV decelerates with the leading vehicle, the velocity of the entire mixed traffic system also begins to oscillate significantly. The following HDVs exhibit significant fluctuations in their headways. Taking the first HDV following behind as an example, the headway drops to 33 meters and then continues to rise to around 72 meters, indicating poor stability due to the promotion of disturbances through the entire system.
@@
Fig. 10 (b) and Fig. 11 (b) shows the mixed traffic system evolution after implementing the LTI-MPC. There is a change that needs to be explained in advance, which is that the first CAV did not maintain a constant velocity with the leading vehicle at the beginning. This is because LTI-MPC imposes a penalty on the error between the actual headway and the expected headway. Once there is a small error between the initial headway set in the experiment and the expected headway after traffic system operates, it will lead to fluctuations in the initial stage of mixed traffic system evolution, which is difficult to be avoided in the experimental setup. It is obvious that the amplitude of the velocity oscillation of the first CAV also begins to oscillate after receiving the oscillation propagation from the leading vehicle, and the following HDV also oscillates with it. The overall oscillation amplitude decreases relative to Basic-MPC. In terms of headway, taking the second HDV indicated by the red line as an example, the oscillation amplitude of the headway improved due to the influence of the CAV ahead, reaching a maximum front distance of about 50 meters at 40 seconds. Subsequently, the second CAV also optimized its driving behavior under the action of LTI-MPC, and the trajectories of the following two trucks also fluctuated accordingly, maintaining a distance near the expected headway.
@@
The evolution and improvement of the mixed traffic system under the Deep-LCC control scenario are shown in Fig. 10 (c) and Fig. 11 (c). Similarly, the initial evolution stage, like the LTI-MPC, requires the first CAV to accelerate in order to achieve equilibrium headway. We can clearly observe that there is little difference between DeeP-LCC and LTI-MPC in overall traffic system evolution improvement, achieving performance comparable to LTI-MPC. The specific details will not be repeated in the description. Of course, Deep-LCC does not require prior knowledge of system dynamics, and its advantage lies in its data-driven non parametric strategy.
@@
Overall, these comparative results further demonstrate that AdapKoopPC can effectively and proactively guide the subsequent HDVs and maintains the smooth driving of the mixed traffic system. More importantly, whether it is LTI-MPC or Deep-LCC, it is necessary to utilize a given model or simulation to generate some trajectories. However, these trajectories are highly dependent on the expected headway, which has limitations. AdapKoopPC adopts a penalty of velocity difference, which matches the expected goal of most drivers of vehicles in a following state. This will not cause discomfort to the driver and is more conducive to optimize the control of the mixed traffic flow system.
6.2.2 Computation time comparison
Table 4 states the real-time computing time for different control methods at each time step, with a time step of 0.12 seconds. The average computation time for LTI-MPC, Deep LCC, and AdadKoopPC is 0.004s, 0.326s, and 0.009s, respectively. To investigate its reasons, LTI-MPC does not require neural network encoding, and the mixed traffic system state dimension is low, so the solution is the fastest. As for Deep-LCC, it was shown to be 0.05s in theprevious work. The online computation time of Deep-LCC has shown significant fluctuations in our experiments. When simulating mixed traffic flow system scenarios similar to pre-generated trajectory scenarios, the average computation delay is 0.326 seconds. When there is a significant difference between the two scenarios, the model training solution needs to iterate to the maximum number of manually set iterations. Therefore, the delay depends on this value and will not be lower than 0.326 seconds, which may be due to a lack of generalization ability. After the impact of neural network encoding time and significant increase in state dimension, the online computation of AdadKoopPC yields an average time of 0.009 seconds, and the sampling interval is 0.12 seconds, which demonstrates the feasibility of implementing real-time mixed traffic system control. It is worth mentioning that the simulation experiments of LTI-MPC and Deep-LCC are based on the MATLAB quadprog optimizer, which has better solving performance compared to the SciPy adopted in AdadKoopPC. Therefore, the online computation time of AdaptKoopPC can still be further improved.